OpenStudy (anonymous):

o.k need help checking my math. Graphing....y=(-3/4)x+5 ...I put -1 and -2 for X. Think I got the others right just checking my work

OpenStudy (anonymous):



OpenStudy (anonymous):

not very good

jimthompson5910 (jim_thompson5910):

\[\large y=-\frac{3}{4}x+5\] \[\large y=-\frac{3}{4}(-1)+5\] \[\large y=\frac{3}{4}+5\] \[\large y=\frac{3}{4}+\frac{20}{4}\] \[\large y=\frac{23}{4}\] So when x = -1, y = 23/4 ----------------------------------------------- \[\large y=-\frac{3}{4}x+5\] \[\large y=-\frac{3}{4}(-2)+5\] \[\large y=\frac{6}{4}+5\] \[\large y=\frac{6}{4}+\frac{20}{4}\] \[\large y=\frac{26}{4}\] So when x = -2, y = 26/4

OpenStudy (anonymous):

so do I divide 26/4 to brake down to where i could plot the points?

jimthompson5910 (jim_thompson5910):

you can, so 26/4 = 6 1/2 = 6.5

OpenStudy (anonymous):

and 23/4 = 5.75?

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

okay i got it a different way.....I turned 3/4 into -.75

OpenStudy (anonymous):

is this okay?

jimthompson5910 (jim_thompson5910):

3/4 = 0.75 not -0.75

OpenStudy (anonymous):

but the it's (-3/4) in the equation...so i turned it into -0.75

jimthompson5910 (jim_thompson5910):

oh ok, forgot about the negative as all I saw was 3/4 So yes, -3/4 = -0.75